The force exerted by the wind on the sails of a sailboat is Fsail = 400 N north. The water exerts a force of Fkeel = 190 N east. If the boat (including its crew) has a mass of 260 kg, what are the magnitude and direction of its acceleration?
magnitude m/s2
direction ° north of east
1
Expert's answer
2019-10-02T10:02:03-0400

Establishing reference system with unit vectors.


  • North = + j
  • East = + i
  • West = -i
  • South = -j

Expressing as a vector each force acting on the boat.


  • Force on the sail

"\vec{F_{Sail}}=400(j)Nw"

  • Forces exerted by water:
  • "\vec{F_{Keel}}=190(i)Nw"


The net Force is equal to:"\vec{F_{Net}}=(190i+400j)Nw"


Applying Newton's second Law on the sailboat


"\vec{F_{net}}=m*\vec{a}"


Where

  • Net force "\vec{F_{Net}}=(190i+400j)Nw"
  • Mass "m=260kg"
  • Acceleration "\vec{a}"


Expression for acceleration:


"\vec{F_{net}}=m*\vec{a} \\ \vec{a}=\frac{\vec{F_{Net}}}{m}"


Evaluating numerically.


"\vec{a}=\frac{\vec{F_{Net}}=(190i+400j)Nw}{260kg} \\\vec{a}=(0.731i+1.54j)\frac{m}{s^{2}}"


The magnitude is given by: "|\vec{a}|=\sqrt{(0.731)^{2}+(1.54)^{2}} \\ |\vec{a}|=1.70\frac{m}{s^{2}}"


The magnitude of the acceleration is:"\boxed{|\vec{a}|=1.70\frac{m}{s^{2}}}"

The angle is given by

: "tan\theta=(\frac{ 1.54}{ 0.731})\\ \theta=tan^{-1}(\frac{ 1.54}{ 0.731}) \\ \theta=64.6^{0}" Expressing respect to the East


The direction towards the east of the acceleration is: "\boxed{64.6^{0}\text{ north of east}}"